OpenStudy (anonymous):

Solve the system of equations. y=2x y=12-x Explain step by step. I'm clueless

OpenStudy (anonymous):

y=2x y=12-x 2x=12-x 3x=12 x=4 Substitution method can you tell me what y is hint(y=2(4)

OpenStudy (anonymous):

I'm thinking 8 ?

OpenStudy (anonymous):

@Ninagf99 ^